A golf ball (mass=1.60oz.) at rest is struck with a golf club ( mass= 8.82oz.) that launches the ball at an angle of 40 degrees. If the initial speed of the golf club is 25.0 m/s and the finial speed of the club is 18.0 m/s, determine the horizontal distance (range) that the golf ball will travel.
Answer #1

let v is the speed of golf ball after it is hit by glof club.

Apply conservation of momentum

8.82*25 = 8.82*18 + 1.6*v

==> v = (8.82*25 - 8.82*18)/1.6

= 38.6 m/s

Horizontal range of golf ball,

R = v^2*sin(2*theta)/g

= 38.6^2*sin(2*40)/9.8

= 149.73 m <<<<<<<<<<<<<<<------------Answer